Mr. John starts his car from home at rest and accelerates uniformly at a rate of 2m/s2 for 2s. He maintains a constant velocity for 10 seconds as he drives to a shopping mall. He then applies the brakes and then the car retards uniformly coming to rest in 5 seconds at the mall. Assuming that the journey is linear, for Mr. John’s car,

(a) draw the motion graph for the whole journey.

(b) find the maximum velocity.

(c) find the retardation in the last 5 seconds.

(d) determine the total distance traveled from home to the shopping mall.

(e) what was the average velocity?



1
Expert's answer
2021-02-28T06:21:10-0500

a = 2m/s2, d = at2/2 = t2 m for the first 2 seconds. d (0s) = 0 m, d (2s) = 4 m.

(b) v = a*t = 2m/s2*2s = 4 m/s, d = d(2s) + V*(t-2) = 4*t m from 2 to 12 seconds. d (12s) = 4 + 4*(12-2) = 44 m.

(c) a = (0 - 4)/5 = -0.8 m/s2, (d) total distance = 44 + 0.8*(17-12)2/2 = 54 m

d = 54-0.8*(x-17)2/2 from 12 to 17 seconds.

(e) average velocity = (54 m)/(17 s) = 3.176 m/s

(a)